# Theoretical investigations of tetrameric magnetic molecules for sub-kelvin cooling

### v1

@@ Abstract
Magnetic molecules are a class of compounds that is also investigated in view of their magnetocaloric properties. The isothermal entropy change and the adiabatic temperature change are key figures of merit for magnetocaloric performance. Here, we investigate spin systems of realistic molecular structures comprising four spins. In view of possible large spin quantum numbers as for gadolinium we model these spin systems by a combination of Heisenberg and dipolar spin-spin interactions. It turns out that a tetrahedral structure with ferromagnetic exchange interactions yields the best figures of merit.
@@ I Introduction
Such an energy gap occurs for instance in spin systems termed single-molecule magnets that are characterized by a strong easy-axis anisotropy that separates a two-fold degenerate ground state from higher lying levels. However, such spin systems show a rotational magnetocaloric effect that occurs when the magnitude of the magnetic field is kept constant and the sample (or the field) are rotated [11, 12, 13, 14]. Although very appealing, technical difficulties such as heat from friction prevent a use at low temperatures so far.
@@
In this article, we are going to investigate spin systems composed of four spins with realistic geometries as obtained in chemical synthesis. This investigation is along the ideas of Refs. [23, 24, 25] where square-pyramidal structures have been studied employing the Heisenberg model. In our current study, we add intramolecular dipolar interaction which can be sizable for larger spins [26]. It also plays an important role below about one kelvin.
Our investigations suggest that tetrahedral arrangements of spins with ferromagnetic exchange interactions yield the strongest magnetocaloric effect. Furthermore, among the investigated structures it is the only arrangement that is able to achieve very low temperatures. We will substantiate our findings with a comparison of four typical structures, however, limit our self to a few central graphs. The reader is encouraged to have a look at a recent Ph.D. thesis offering detailed explorations of the possible huge parameter space [27].
@@ II Model
The spin systems are treated by exact diagonalization. Thermal equilibrium observables are calculated in the canonical ensemble. We concentrate on realistic exchange interactions and a realistic field change from $B=7$ T to $B=0$ when evaluating the isothermal entropy change.
In the following we consider four different typical spin systems as listed in table 1, a tetrahedron, a butterfly system, a linear chain, and a square. We allow two different exchange interactions in our search for advantageous coupling schemes.
@@ III Results
In the following we discuss four typical motives as they appear in molecular compounds. We show isothermal entropy changes in the parameter space of $J_{1}$ and $J_{2}$ for the spin quantum number $s=3/2$ for a field sweep from $B=7$ T down to $B=0$. The Heisenberg limit is denoted as $d=\infty$, i.e., for zero dipolar interaction. In order to account for dipolar influence we picked some reasonable (but not too large) distances. It is clear that the parameter space of such an investigation is practically limitless since one could vary the spin quantum numbers, introduce more exchange constants, and vary the distances that are important for the dipolar interaction in a much larger range.
Here we focus on central and typical results, the reader is invited to have a look at the wealth of simulation results presented in the Ph.D. thesis [27].
Figure 1: Isothermal entropy change for a tetrahedron with $s=3/2$ at $T=10$ K (top), $T=0.1$ K (middle), and $T=0.1$ K (bottom); the latter including dipolar interactions for a distance of $d=2$ Å. The magnetic field is perpendicular to the plane of the red triangle in the tetrahedron, see Tab. 1. The color maps show the respective entropy changes.
For the discussion of the isothermal entropy change we picked two temperatures, $T=10$ K and $T=0.1$ K, again more in [27]. The behavior at $T=0.1$ K without dipolar interactions is then contrasted with the behavior including dipolar interactions. At $T=10$ K the dipolar interaction has practically no effect.
Figure 2: Isothermal entropy change for a butterfly with $s=3/2$ at $T=10$ K (top), $T=0.1$ K (middle), and $T=0.1$ K (bottom); the latter including dipolar interactions for distances of $d_{1}=2.3$ Å (corresponding to $J_{13}$) and $d_{2}=3.3$ Å (corresponding to $J_{24}$). The magnetic field points along the longer diagonal of the butterfly corresponding to $J_{24}$, see Tab. 1. The color maps show the respective entropy changes.
@@
Figure 3: Isothermal entropy change for a chain with $s=3/2$ at $T=10$ K (top), $T=0.1$ K (middle), and $T=0.1$ K (bottom); the latter including dipolar interactions for distances of $d=3$ Å. The magnetic field points along the axis of the chain. The color maps show the respective entropy changes.
@@
The butterfly and the chain, Figs. 2 and 3, loose their good isothermal entropy changes for ferromagnetic couplings at sub-kelvin temperatures when the dipolar interaction is taken into account. This effect is rather drastic as can be seen in the bottom panels of both figures. For $T=10$ K, the chain shows good isothermal entropy changes for small ferromagnetic interactions whereas the butterfly performs good for ferromagnetic $J_{\text{wings}}$ independent of the sign of $J_{\text{body}}$, compare top panel of Fig. 2. In view of existing chemical compound this is a clear advantage.
@@ IV Discussion
In this paper we investigated four typical geometric structures of four spins each as they often appear as motives in magnetic molecules. It turns out that a tetrahedral arrangement with ferromagnetic exchange interactions is most promising in view of the isothermal entropy change that is robust under the influence of dipolar interactions at sub-kelvin temperatures. Even combinations of ferromagnetic and antiferromagnetic interactions yield a reasonable magnetocaloric performance for the tetrahedron.
@@
Finally, Fig. 6 compares the four investigated tetranuclear spin systems with ferromagnetic coupling (J=20J=-20 K) in view of the adiabatic temperature change. Again, $B_{\text{hot}}=7$ T, $B_{\text{cold}}=0$, and $T_{\text{hot}}=10$ K. One sees that the tetrahedron outperforms all other three structures in terms of the achievable $T_{\text{cold}}$ for every investigated distance of spins, i.e., for every strength of the dipolar interaction. This seems to be largely independent of spin quantum number. We speculate that in the case of the tetrahedron the dipolar interaction does not split the ground multiplet as much as it does for the other three structures. The magnitude of this splitting sets the scale for the lowest achievable temperatures.
However and unfortunately, tetrahedra with larger spins – iron or gadolinium – seem to dominantly exhibit antiferromagnetic interactions, see e.g. [30, 31, 32, 33, 34, 35]. Nickel-based tetrahedra may show ferromagnetic exchanges, see e.g. [36, 37] as well as some rare manganese compounds [38] do, although for nickel also antiferromagnetic couplings are reported [39]. However, the larger problem with nickel and manganese is the non-negligible single-ion anisotropy and the possible biquadratic exchange for nickel that would very likely alter the magnetocaloric behavior at sub-kelvin temperatures.

### v2

@@ Abstract
Magnetic molecules are a class of compounds that is also investigated in view of their magnetocaloric properties. The isothermal entropy change and the adiabatic temperature change are key figures of merit for magnetocaloric performance. Here, we investigate spin systems of realistic molecular structures comprising four spins. In view of potentially large spin quantum numbers as for gadolinium we model these spin systems by a combination of Heisenberg and dipolar spin-spin interactions. It turns out that a tetrahedral structure with ferromagnetic exchange interactions yields the best figures of merit.
@@ I Introduction
Such an energy gap occurs for instance in spin systems termed single-molecule magnets that are characterized by a strong easy-axis anisotropy that separates a two-fold degenerate ground state from higher-lying levels. However, such spin systems show a rotational magnetocaloric effect that occurs when the magnitude of the magnetic field is kept constant and the sample (or the field) are rotated [11, 12, 13, 14]. Although very appealing, technical difficulties such as heat from friction prevent a use at low temperatures so far.
@@
In this article, we are going to investigate spin systems composed of four spins with realistic geometries as obtained in chemical synthesis. This investigation is along the ideas of Refs. [23, 24, 25] where square-pyramidal structures have been studied employing the Heisenberg model. In our current study, we add intramolecular dipolar interaction which can be sizable for larger spins [26]. It also plays an important role below about one kelvin. For reasons of computational feasibility we study only single-spin quantum numbers of $s=3/2$ in this article. The qualitative conclusions are not altered by this restriction.
Our investigations suggest that tetrahedral arrangements of spins with ferromagnetic exchange interactions yield the strongest magnetocaloric effect. Furthermore, among the investigated structures it is the only arrangement that is able to achieve very low temperatures. We will substantiate our findings with a comparison of four typical geometrical structures, however, limit ourselves to a few central graphs. The reader is encouraged to have a look at a recent Ph.D. thesis (in German) offering detailed explorations of the possible huge parameter space [27].
@@ II Model
The spin systems are treated by numerically diagonalizing the Hamiltonian matrix for every investigated pair of exchange interactions, every external field and every dipolar arrangement. Thermal equilibrium observables are calculated in the canonical ensemble using the eigenvalues and the eigenvectors of the Hamiltonian matrix. We concentrate on realistic exchange interactions and a realistic field change from $B=7$ T to $B=0$ when evaluating the isothermal entropy change.
@@ III Results
In the following, we consider four different typical spin systems as listed in table 1, a tetrahedron, a butterfly system, a linear chain, and a square. We allow two different exchange interactions in our search for advantageous coupling schemes. We show isothermal entropy changes in the parameter space of $J_{1}$ and $J_{2}$ for the spin quantum number $s=3/2$ for a field sweep from $B=7$ T down to $B=0$. The Heisenberg limit is denoted as $d=\infty$, i.e., for zero dipolar interaction. In order to account for dipolar influence we pick some reasonable (but not too large) distances. It is clear that the parameter space of such an investigation is practically limitless since one could also vary the spin quantum numbers, introduce more exchange constants, and vary the distances that are important for the dipolar interaction in a much larger range. Here we focus on central and typical results, see also [27].
Figure 1: Isothermal entropy change for a tetrahedron with $s=3/2$ at $T=10$ K (top), $T=0.1$ K (middle), and $T=0.1$ K (bottom); the latter including dipolar interactions for a distance between of $d=2$ Å. The magnetic field is perpendicular to the plane of the red triangle in the tetrahedron, see Tab. 1. The color maps show the respective entropy changes.
For the discussion of the isothermal entropy change we picked two temperatures, $T=10$ K and $T=0.1$ K. The behavior at $T=0.1$ K without dipolar interactions is then contrasted with the behavior including dipolar interactions. At $T=10$ K the dipolar interaction has practically no effect.
Figure 2: Isothermal entropy change for a butterfly with $s=3/2$ at $T=10$ K (top), $T=0.1$ K (middle), and $T=0.1$ K (bottom); the latter including dipolar interactions for distances of $d_{1}=2.3$ Å (corresponding to the exchange path $J_{13}$) and $d_{2}=3.3$ Å (corresponding to the exchange path $J_{24}$). The magnetic field points along the longer diagonal of the butterfly corresponding to the exchange path $J_{24}$, see Tab. 1. The color maps show the respective entropy changes.
@@
Figure 3: Isothermal entropy change for a chain with $s=3/2$ at $T=10$ K (top), $T=0.1$ K (middle), and $T=0.1$ K (bottom); the latter including dipolar interactions for distances between adjacent spins of $d=3$ Å. The magnetic field points along the axis of the chain. The color maps show the respective entropy changes.
@@
The butterfly and the chain, Figs. 2 and 3, lose their good isothermal entropy changes for ferromagnetic couplings at sub-kelvin temperatures when the dipolar interaction is taken into account. This effect is rather drastic as can be seen in the bottom panels of both figures. For $T=10$ K, the chain shows good isothermal entropy changes for small ferromagnetic interactions whereas the butterfly performs good for ferromagnetic $J_{\text{wings}}$ independent of the sign of $J_{\text{body}}$, compare top panel of Fig. 2. In view of existing chemical compound this is a clear advantage.
@@ IV Discussion
In this paper, we investigated four typical geometric structures of four spins each as they often appear as motives in magnetic molecules. It turns out that a tetrahedral arrangement with ferromagnetic exchange interactions is most promising in view of the isothermal entropy change that is robust under the influence of dipolar interactions at sub-kelvin temperatures. Even combinations of ferromagnetic and antiferromagnetic interactions yield a reasonable magnetocaloric performance for the tetrahedron.
@@
Finally, Fig. 6 compares the four investigated tetranuclear spin systems with ferromagnetic coupling (J=20J=-20 K) in view of the adiabatic temperature change.
Again, $B_{\text{hot}}=7$ T, $B_{\text{cold}}=0$, and $T_{\text{hot}}=10$ K. One sees that the tetrahedron outperforms all other three structures in terms of the achievable $T_{\text{cold}}$ for every investigated distance of spins, i.e., for every strength of the dipolar interaction. This seems to be largely independent of spin quantum number. We speculate that in the case of the tetrahedron the dipolar interaction does not split the ground multiplet as much as it does for the other three structures. The magnitude of this splitting sets the scale for the lowest achievable temperatures.
However and unfortunately, tetrahedra with larger spins – iron(III) or gadolinium(III) – seem to dominantly exhibit antiferromagnetic interactions, see e.g. [30, 31, 32, 33, 34, 35].
@@
Nickel-based tetrahedra may show ferromagnetic exchanges, see e.g. [36, 37] as well as some rare manganese compounds [38] do, although for nickel also antiferromagnetic couplings are reported [39]. However, the larger problem with nickel and manganese is the non-negligible single-ion anisotropy and the possible biquadratic exchange for nickel that would very likely alter the magnetocaloric behavior at sub-kelvin temperatures in an unfavorable way [1].